Visiuni Useu Find explicit formulas for sequences of the form az, az, az, ... with the initial terms given below.
Answer #1

= +1, Q2 = 6+Az = 4 + ſi By observing the above terms, we get idea that Ian = 4th + 1 - MEN Scanned with CamScanner